OpenStudy (anonymous):

How do I solve dy/dx = (y-1)^2? y(0)=1

OpenStudy (anonymous):

dy/(y-1)^2 = dx -1/(y-1) = x+c

OpenStudy (anonymous):

yd/dx (y-1 ^ 2 ) = (y-1)^(2-1) y(0) = 0-1 = -1 !

OpenStudy (anonymous):

this must be wrong///

OpenStudy (anonymous):

it must be dx/dy instead of dy/dx

OpenStudy (anonymous):

my teacher mentioned this type of problem

OpenStudy (anonymous):

-1/(x+c) = y-1 y=(x+c-1)/(x+c) now put the values 0,1 c-1/c = 0 c=1

OpenStudy (anonymous):

there you go

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

click on good answer to complete my century plz ;)

OpenStudy (anonymous):

thnx